OpenStudy (anonymous):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. C = 37°, a = 19, c = 8

OpenStudy (campbell_st):

so the law of sines says \[\frac{\sin(A)}{a} = \frac{\sin(C)}{c}\] then using your information \[\sin(A) = \frac{19 \times \sin(C)}{8}\] for a triangle to be formed from the information -1 <= sin(A) <= 1 so just find the value of sin(A) using the information above...